OpenStudy (yanasidlinskiy):

@jim_thompson5910 Do you mind helping me with quadratic Formulas?

OpenStudy (yanasidlinskiy):

Like, I wanna know how to do them.

OpenStudy (puzzler7):

For a quadratic equation \[ax^2+bx+c=0\]\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (puzzler7):

Just plug your a, b, and c into the equation to find x.

OpenStudy (yanasidlinskiy):

Ok..like..Can you give me an example?

jimthompson5910 (jim_thompson5910):

say you want to solve 2x^2 - 3x - 20 = 0 for x

jimthompson5910 (jim_thompson5910):

compare 2x^2 - 3x - 20 to ax^2 + bx + c we see that a = 2 b = -3 c = -20

jimthompson5910 (jim_thompson5910):

plug these values into the quadratic formula \[\large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] \[\large x=\frac{-(-3) \pm \sqrt{(-3)^2-4(2)(-20)}}{2(2)}\] \[\large x=\frac{3 \pm \sqrt{169}}{4}\] \[\large x=\frac{3 \pm 13}{4}\] \[\large x=\frac{3 + 13}{4} \ \text{ or } \ x = \frac{3 - 13}{4}\] \[\large x=\frac{16}{4} \ \text{ or } \ x = \frac{-10}{4}\] \[\large x=4 \ \text{ or } \ x = -\frac{5}{2}\] -------------------------------------------------------------------- So the two solutions to 2x^2 - 3x - 20 = 0 are x = 4 or x = -5/2

OpenStudy (yanasidlinskiy):

Ok. I *think I got it. What If i had the problem y^2-81=0.

OpenStudy (yanasidlinskiy):

What would I do then?

jimthompson5910 (jim_thompson5910):

if you want to use the quadratic formula, then think of y^2-81 as y^2+0y-81

jimthompson5910 (jim_thompson5910):

so a = 1 b = 0 c = -81

OpenStudy (yanasidlinskiy):

Oh...Ok...I need help on this though.. x^2 - 3x - 10 = 0...Im going to try it and see what I get. Can you tell me if I got it right or wrong?

jimthompson5910 (jim_thompson5910):

sure, post what you get

OpenStudy (yanasidlinskiy):

My: a:1 b:-3 c:-10. Right?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

plug those values into the quadratic formula

OpenStudy (yanasidlinskiy):

I got: \[x=\frac{ 3\pm \sqrt{-50} }{ 2 }\] Is that right?

jimthompson5910 (jim_thompson5910):

b^2 - 4ac = (-3)^2 - 4(1)(-10) b^2 - 4ac = 9 + 40 b^2 - 4ac = 49

jimthompson5910 (jim_thompson5910):

not sure how you got -50

OpenStudy (yanasidlinskiy):

I got it by multiplying the - times a positive..Like i did it in order. Multiplied the 3 subtracted that from 4 then multiplied it by 1 then multiplied it by -10. I dunno..Lemme check

jimthompson5910 (jim_thompson5910):

do you see how I got 49?

OpenStudy (yanasidlinskiy):

Ooooo...Ok. I got it. I did it the wrong way....I see what you did. What would I do after that step?

jimthompson5910 (jim_thompson5910):

you break up the plus/minus to get these two equations x = (3 + 49)/2 or x = (3 - 49)/2

jimthompson5910 (jim_thompson5910):

oops sry

OpenStudy (yanasidlinskiy):

I think it'd be: \[x= \frac{ 3\pm \sqrt{7} }{ 2 }\]

jimthompson5910 (jim_thompson5910):

the square root of 49 is 7, so these two equations x = (3 + 7)/2 or x = (3 - 7)/2

jimthompson5910 (jim_thompson5910):

once you take the square root of a number, the square root symbol goes away

OpenStudy (yanasidlinskiy):

oh. Ok. So that way it'll be just + or -. Got it!!!!:) Thank You!!!!

jimthompson5910 (jim_thompson5910):

you're welcome